OpenStudy (anonymous):

Medals and Fan Given What is the quotient of nine over ten divided by six over eleven in simplest form? one and thirteen over twenty one and one over ten twenty-seven over fifty-five fifteen over twenty-one

OpenStudy (anonymous):

(9/10)÷(6/11) means (9/10)*(11/6) So we have ((9*11)/(10*6)) Or simply 33/20 Which is 1 and 13/20

OpenStudy (texaschic101):

(9/10) / (6/11) To divide fractions, " flip " what you are dividing by, then multiply. (9/10) / (6/11) = 9/10 * 11/6 = 99/60 = 33/20 or 1 13/20 Be sure not to flip the other number... (9/10) / (6/11)...if you did this..10/9 * 6/11 (flipped the other number). then it would be wrong.

OpenStudy (anonymous):

Thanks that helped me a lot. That was extremely help full. :)

OpenStudy (texaschic101):

glad to help :)